A balloon having 8gm Ne gas and radius 20 cm is pricked and 7gm of Ne gas effused from it. What will be the radius of the balloon under similar condition

Grade:11

2 Answers

Himanshu Angane
4 years ago
Hi friend 
PV=nRT
For neon
P×4/3×3.14×20^3=8/20×R×T
You'll get
83733=RT/P
After effused 7 gram reaming 1 gram
PV=nRT
PC=1/20×RT
20V=RT/P
FROM ABOVE 
20V=83733
V=83733/20
V=4186
Now 
R^3=4168×3/3.14×4
R^3=1000
R=10cm
Is the answer
 
83733=after effused 7 gram
 
shethsaumya03
4 years ago
PV = nRT